# An empirical appraisal of methods for the dynamic prediction of survival with numerous longitudinal predictors

### v1

@@ Abstract
Recently, the increasing availability of repeated measurements in biomedical studies has motivated the development of several statistical methods for the dynamic prediction of survival in settings where a large (potentially high-dimensional) number of longitudinal covariates is available. These methods differ in both how they model the longitudinal covariates trajectories, and how they specify the relationship between the longitudinal covariates and the survival outcome. Because these methods are still quite new, little is known about their applicability, limitations and performance when applied to real-world data.
To investigate these questions, we present a comparison of the predictive performance of the aforementioned methods and two simpler prediction approaches to three datasets that differ in terms of outcome type, sample size, number of longitudinal covariates and length of follow-up. We discuss how different modelling choices can have an impact on the possibility to accommodate unbalanced study designs and on computing time, and compare the predictive performance of the different approaches using a range of performance measures and landmark times.
Keywords: dynamic prediction; survival analysis; longitudinal data; risk prediction modelling.
@@ 1 Introduction
Estimating the probability that individuals may experience adverse events (e.g., onset of a disease or death) is an important task of modern medicine. Risk prediction models (Steyerberg,, 2009) are statistical and machine learning models that can be employed to estimate the probability that each individual will experience a certain event over time. Risk prediction models can be used by clinicians to monitor disease progression, provide patients with information about their health status and risks, and guide decisions about hospitalization, surgeries and initiating, adjusting or terminating a certain treatment.
@@
A common approach to risk prediction modelling involves estimating the probability that individuals will not experience the event over time (called survival probability) based on a set of risk factors (covariates) that are either time-independent, i.e. do not change over time or are measured at a single time point. The main advantages of this static modelling approach are that it is relatively easy to implement and it does not require gathering patient information over time. However, predictions developed in this way do not incorporate any information on individual changes over time that may be predictive of the event of interest. If such information is deemed relevant for the occurrence of the event, then ideally we would like to measure time-varying covariates at multiple points in time, and to incorporate information from such repeated measurements in the prediction model to improve predictive accuracy. Moreover, we would like to be able to dynamically update individual predictions of survival each time new (more recent) measurements become available.
@@
Traditional approaches to the dynamic prediction of survival include landmarking (Van Houwelingen,, 2007) and joint modelling (Rizopoulos,, 2012). Landmarking requires the specification of a so-called landmark time. Individuals that experienced the event or were censored before the landmark are discarded, so that modelling is restricted to individuals who survived up to the landmark time. Then, for each longitudinal covariate a summary measure is computed from all repeated measurements up to the landmark time, and the summary measures thus computed are used as predictors in a Cox proportional hazards model. In the simplest approach, called last observation carried forward (LOCF) landmarking, the last observation of each longitudinal covariate before the landmark time is used as summary measure. The major advantages of the LOFC approach are its simplicity and ease of implementation; possible disadvantages include the fact that the last observation taken before the landmark may be outdated, and the lack of a model to describe the evolution over time of the longitudinal covariates and to account for measurement error (which may be considerable in the case of biomarkers). Joint modelling involves the specification and estimation of a shared random effects model for both the longitudinal covariates and the survival outcomes. Its estimation is typically computationally demanding, and this has so far restricted the applicability of joint models to problems with only one or at most a handful of longitudinal covariates. Recently, Putter and van Houwelingen, (2022) proposed a “landmarking 2.0” approach that tries to combine different aspects of landmarking and JMs, focusing on a situation with only one longitudinal biomarker.
@@
Over the last decades, technological advancements have led to an increasing availability of studies where many longitudinal covariates are measured alongside with a survival outcome. Examples of such studies include observational studies were patients are monitored for several years to study the occurrence of a slowly progressing disease such as cancer (Patel et al.,, 2017), dementia (Bennett et al.,, 2018) or Alzheimer’s disease (Weiner et al.,, 2010), as well as studies focused on the identification of biomarkers associated with disease progression and the occurrence of disease milestones among a large number of longitudinal omic (e.g., protein, gene or metabolite expression levels) markers (Signorelli et al.,, 2020; Filbin et al.,, 2021).
@@
Recently, this increasing availability of repeated measurement data in biomedical studies has motivated the development of new statistical methods that can be used for the dynamic prediction of survival in situations where a large (potentially high-dimensional) number of longitudinal covariates is available. These approaches differ in the way in which they model the evolution over time of the longitudinal covariates, as well as in the way in which they model the relationship between summaries of the longitudinal covariates and the survival outcome. Li and Luo, (2019) proposed the Multivariate Functional Principal Component Analysis Cox model (MFPCCox), which employs multivariate functional principal component analysis (MFPCA) to summarize the longitudinal covariates, and then uses the MFPCA scores as covariates in a Cox model for the survival outcome. Signorelli et al., (2021) proposed Penalized Regression Calibration (PRC), which uses linear as well as multivariate mixed-effects models to model the trajectories of the longitudinal predictors, and predicts the survival outcome with a penalized Cox model where the predicted random effects are included as covariates. Although Signorelli et al., (2021) did not consider a dynamic prediction setup, the use of PRC for dynamic prediction was later illustrated in Signorelli, (2023). Lin et al., (2021) proposed the Functional Random Survival Forest (FunRSF) approach, where longitudinal covariates are modelled through MFPCA similarly to MFPCCox, and the MFPCA scores are then used as time-fixed covariates to predict survival with a random survival forest. Lastly, Devaux et al., (2022) proposed a dynamic Random Survival Forest (DynForest) method that uses linear mixed models (LMMs) to summarize the longitudinal covariates, and then uses the predicted random effects as time-fixed covariates within a random survival forest.
@@
Because these new methods were proposed in the last few years, to date little is known about their applicability to, and performance with, real-world datasets. Practical modelling questions include how well each method can accommodate datasets with unbalanced repeated measurements, and how computationally intensive each method is as a function of the sample size and number of longitudinal covariates (specially when compared to a computationally inexpensive approach such as landmarking and to a computationally inefficient one such as JM). Furthermore, it is of primary interest to evaluate the predictive performance of these alternative modelling approaches on multiple real-world datasets with different features.
@@
To investigate these questions, we gathered data from three longitudinal studies that differ in terms of outcome type, sample size, number of longitudinal covariates and length of follow-up. We applied MFPCCox, PRC, FunSRF and DynForest, alongside two simpler prediction approaches, to each dataset, and proceeded to evaluate the predictive performance of each method according to three different metrics for survival data.
@@
The remainder of this article is organized as follows: in Section 2 we briefly describe the problem of dynamic prediction and the methods considered in our empirical comparison. In Section 3 we present the 3 datasets included in our comparison alongside with information about how the prediction methods were applied to each dataset. Section 4 describes the results of the comparison, whereas Section 5 contains some discussion points and concluding remarks.
2 Methods for the dynamic prediction of survival
2.1 Dynamic prediction of survival
@@
We consider the setup of a longitudinal study where $n$ subjects are enrolled, and interest lies in predicting the occurrence of a certain event of interest for the different individuals. Each subject $i\in\{1,2,...,n\}$ is followed from time $t=0$ (study entry) until they either experience the event of interest at time $t_{i}$, or they are right-censored at time $c_{i}$. We denote by $T_{i}$ and $C_{i}$ the random variables corresponding to the event and censoring times, respectively, and let $T_{i}^{*}=\min{\left(T_{i},C_{i}\right)}$. Furthermore, we denote by $t_{i}^{*}$ the observed value of $T_{i}^{*}$ for subject $i$, and introduce a censoring indicator $\delta_{i}$ that is $1$ if for subject $i$ the event of interest is observed at $t_{i}^{*}=t_{i}$, and 00 in case of censoring, i.e. if the event has not yet been observed at $t_{i}^{*}=c_{i}$.
@@
Between times 0 and $t_{i}^{*}$, subject $i$ undergoes $m_{i}\geq 1$ visits at times $t_{i1},t_{i2},...,t_{im_{i}}$, where we assume that $t_{i1}=0$, $t_{i1}<t_{i2}<\ldots<t_{im_{i}}$ and $t_{im_{i}}<t_{i}^{*}$. At time $t_{i1}=0$, $P$ baseline covariates $x_{i}=(x_{1i},...,x_{Pi})$ are measured. Moreover, $Q$ longitudinal covariates are measured at each follow-up visit (including $t_{i1}=0$). We equivalently denote by $y_{qij}=y_{iq}(t_{ij})$ the value of the $q$-th longitudinal covariate measured on subject $i$ at time $t_{ij}$, and let $y_{ij}=y_{i}(t_{ij})=(y_{1ij},...,y_{Qij})$ be the corresponding Q-dimensional vector.
@@
Let $\mathcal{I}(t)=\{i\in\{1,2,...,n\}:t_{i}^{*}>t\}$ denote the set of individuals that are still at risk at a given time point $t$. Moreover, let $\mathcal{H}_{i}(t)=\left(x_{i},y_{i}(0),...,y_{i}(t_{ik})\right)$, where $k$ is such that $t_{ik}\leq t<t_{i,k+1}$, denote the covariate information available for a subject $i\in\mathcal{I}(t)$ up until time $t$ (in other words, $\mathcal{H}_{i}(t)$ contains both the $P$ baseline covariates, and all repeated measurements taken up until time $t$ of the $Q$ longitudinal covariates).
@@
At any point in time $t_{1}\in(0,t_{i}^{*})$, we may want to predict the conditional probability $P(T_{i}>t_{2}|T_{i}>t_{1})$ that individual $i$ will still be event-free at time $t_{2}$, given that they had not experienced the event at $t_{1}<t_{2}$. More specifically, given a so-called landmark time $\ell$, our goal is to estimate the conditional survival function
@@
In dynamic prediction, it can often be of interest to compute predictions of survival $S_{i}(t|\ell_{k},\mathcal{H}_{i}(\ell_{k}))$ over a range of $K$ landmark times $\ell_{1},\ell_{2},...,\ell_{K}$, rather than for a single landmark $\ell$; if that is indeed the case, our notation can be easily adjusted by re-defining $\mathcal{I}(\ell_{k})$ and $\mathcal{H}_{i}(\ell_{k})$ for each landmark time $\ell_{k}$.
2.2 Strict vs relaxed landmarking of the longitudinal data
Equation (1) states that the goal of dynamic prediction is to predict the probability of survival for those individuals that survived past the landmark time $\ell$, using as predictors all covariate values measured up to time $\ell$.
When estimating a statistical model to address the dynamic prediction problem at time $\ell$, one would usually estimate the prediction model using only those covariate values contained in $\mathcal{H}_{i}(\ell)$, discarding any repeated measurement taken after $\ell$ as such measurements would not be available to compute predictions at the prediction time $\ell$. Gomon et al., (2024) named this approach “strict landmarking”, contrasting it to a “relaxed landmarking” approach where the model is estimated including longitudinal measurements taken after $\ell$ for all subjects in the training set.
In the literature, dynamic prediction methods are often estimated using the relaxed data landmarking approach, probably because such approach allows to use more repeated measurements to model the trajectories of the longitudinal data. However, relaxed data landmarking can be problematic for two reasons: first, it uses future information (collected after $t>\ell$) to predict the conditional survival probability $P(T>t|T>\ell)$. Note that such future information is not available at the landmark time, and thus in reality it would not be possible to make use of it to make predictions at $\ell$. Second, relaxed data landmarking introduces a selection bias in the modelling of the longitudinal trajectories: because individuals who survive longer are likely to have more repeated measurements taken after $\ell$ than individuals with shorter survival, the available measurements after the landmark are not a representative sample from the population of individuals who survived up until $\ell$. Whereas this is not an issue with joint models, the two-step approaches considered in this article do not correct for this source of bias. In practice, this means that even though relaxed data landmarking allows to use more data to train these two-steps prediction models, in practice it introduces selection biases that can severely undermine predictive performance. Gomon et al., (2024) compared strict and relaxed data landmarking for the MFPCCox method, showing how models trained with a strict data landmarking approach have superior predictive performance over models trained with relaxed data landmarking. Based on the aforementioned remarks and the results of Gomon et al., (2024), hereafter we describe and implement the dynamic prediction models included in our comparison using a strict data landmarking approach.
2.3 Methods for dynamic prediction

### v2

# Benchmarking multi-step methods for the dynamic prediction of survival with numerous longitudinal predictors

@@ Abstract
In recent years, the growing availability of biomedical datasets featuring numerous longitudinal covariates has motivated the development of several multi-step methods for the dynamic prediction of survival outcomes. These methods employ either mixed-effects models or multivariate functional principal component analysis to model and summarize the longitudinal covariates’ evolution over time. Then, they use Cox models or random survival forests to predict survival probabilities, using as covariates both baseline variables and the summaries of the longitudinal variables obtained in the previous modelling step.
Because these multi-step methods are still quite new, to date little is known about their applicability, limitations, and predictive performance when applied to real-world data. To gain a better understanding of these aspects, we performed a benchmarking of these multi-step methods (and two simpler prediction approaches) using three datasets that differ in sample size, number of longitudinal covariates and length of follow-up. We discuss the different modelling choices made by these methods, and some adjustments that one may need to do in order to be able to apply them to real-world data. Furthermore, we compare their predictive performance using multiple performance measures and landmark times, assess their computing time, and discuss their strengths and limitations.
Keywords: dynamic prediction; survival analysis; longitudinal data; risk prediction models; clinical prediction models.
@@ 1.1 Dynamic prediction
Predicting the probability that individuals may experience adverse events (such as the onset of a disease, or death) is an important task of modern medicine. Risk prediction models (RPMs, Steyerberg, , 2009) are statistical and machine learning models that can be employed to predict the probability that an individual (or patient) will experience an event of interest over time. They can be used by clinicians to monitor disease progression, provide patients with information about their health status and risks, and guide decisions about hospitalization, surgeries and initiating, adjusting or terminating a certain treatment.
@@
The development of a RPM requires the estimation of the probability that individuals will not experience the event over time (called survival probability) as a function of their risk factors (covariates or predictors). Often, all covariates are time-independent (i.e., they do not change over time) or measured at a single point in time. The main advantages of this static modelling approach are that it is relatively easy to implement, and it does not require gathering patient information over time. However, static RPMs do not incorporate any information on individual changes over time that may be predictive of the event of interest. Dynamic RPMs make it possible to include longitudinal (“time-dependent”) covariates (i.e., covariates measured at multiple points in time on the same individual), empowering more accurate predictions in situations where changes in the level of one or more covariates over time are associated with the risk of experiencing the event. Moreover, dynamic RPMs can be used to dynamically update individual predictions of survival each time new (more recent) measurements become available.
1.2 Traditional dynamic prediction methods
Traditional approaches to the dynamic prediction of survival outcomes include landmarking (Van Houwelingen, , 2007) and joint modelling (Hickey et al., , 2016).
@@ 1.2 Traditional dynamic prediction methods
The landmarking method specifies a conditional model for the probability of experiencing the event in the future, given that an individual has not yet experienced the event up until a prespecified landmark time. Individuals that either experienced the event or were censored before the landmark are discarded, while individuals who are still at risk at the landmark time are retained for analysis. Then, all repeated measurements of each longitudinal covariate taken up until the landmark time are summarized into a single summary measure, and the summary measures of the longitudinal covariates are used as predictors in a Cox proportional hazards model (Cox, , 1972) alongside the baseline covariates. In the simplest and most commonly used approach, called last observation carried forward (LOCF) landmarking, the last observation of each longitudinal covariate before the landmark time is used as summary measure. The major advantages of the LOFC approach are its simplicity and ease of implementation; possible disadvantages include the fact that the last observation taken before the landmark may be outdated, and the lack of a model to describe the evolution over time of the longitudinal covariates and to account for measurement error (which may be considerable in the case of biomarkers).
Joint modelling involves the specification and estimation of a shared random effects model (called “joint model”) for both the longitudinal covariates and the survival outcomes. Usually, the joint model (JM) assumes the survival outcome and the longitudinal covariates to be conditionally independent given a vector of subject-specific random effects. Unlike LOCF landmarking, JMs enable efficient use of the longitudinal data, and to account for measurement error. However, their estimation is computationally difficult and demanding, particularly when attempting to include many longitudinal covariates in the joint model. To date, high computing times and frequent convergence problems have limited the applicability of JMs to problems with a limited number of longitudinal covariates (Fournier et al., , 2019; Spreafico and Ieva, , 2021), often forcing researchers working on data with numerous longitudinal covariates to preselect a subset of those covariates before fitting the JM.
1.3 Multi-step dynamic prediction methods
@@
Over the last decades, technological advancements have led to a growing availability of studies where many longitudinal covariates are measured alongside a survival outcome. Examples include observational studies were patients are monitored for several years to study the occurrence of a slowly progressing disease such as cancer (Patel et al., , 2017), dementia (Bennett et al., , 2018) or Alzheimer’s disease (Weiner et al., , 2010), as well as studies that seek to identify biomarkers associated with disease progression and the occurrence of disease milestones among a large number of longitudinal omic variables (Signorelli et al., , 2020; Filbin et al., , 2021).
@@
Recently, this increasing availability of repeated measurement data in biomedical studies has driven the development of new statistical methods (Li and Luo, , 2019; Signorelli et al., , 2021; Lin et al., , 2021; Devaux et al., , 2023; Gomon et al., , 2024) and software (Signorelli, , 2024; Devaux et al., , 2024) for dynamic prediction that can better handle the availability of a large number of longitudinal covariates. These methods approach the dynamic prediction problem through a multi-step modelling approach: first, they employ either mixed-effects models (MEMs) or multivariate functional principal component analysis (MFPCA) to describe the evolution over time of the longitudinal covariates and to obtain time-independent summaries of the repeated measurement data; then, they model the relationship between the survival outcome and the available covariates using either Cox models or random survival forests (RSFs).
In particular, the Multivariate Functional Principal Component Analysis Cox model (MFPCCox, Li and Luo, (2019); Gomon et al., (2024)) uses MFPCA to reduce the longitudinal covariates into subject-specific MFPCA scores, and then proceeds to estimate a Cox model where both baseline covariates and the MFPCA scores are included as covariates. Penalized Regression Calibration (PRC, Signorelli et al., (2021); Signorelli, (2024)) makes use of linear mixed models (LMMs) as well as multivariate mixed models to reduce the longitudinal covariates into subject-specific random effects, then it employs a penalized Cox model to predict the survival outcome based on the predicted random effects and baseline covariates. Similarly to MFPCCox, the Functional Random Survival Forest (FunRSF, Lin et al., (2021)) uses MFPCA to obtain MFPCA scores, but then it uses a random survival forest to predict survival based on the MFPCA scores and baseline covariates. Lastly, similarly to PRC, the dynamic Random Survival Forest (DynForest, Devaux et al., (2023, 2024)) method uses LMMs to compute the predicted random effects; however, it uses the predicted random effects and baseline covariates to predict survival through a random survival forest.
1.4 Benchmarking multi-step dynamic prediction methods
@@
Because these multi-step methods have been introduced very recently, to date little is known about their applicability to, and performance with, real-world datasets. Practical modelling questions include how well these methods can accommodate real-world datasets with numerous longitudinal covariates and unbalanced measurement times, and how computationally intensive they are (especially when compared to a computationally inexpensive approach such as landmarking, and to a computationally inefficient one such as JM). From an applied perspective, it is also unknown what the predictive performance of the different multi-step methods may be when applied to real-world datasets.
@@
The goal of this study is to investigate these open questions using publicly available real-world datasets. We gathered data from three longitudinal studies that differ substantially in terms of sample size, number of longitudinal covariates and length of follow-up. We used MFPCCox, PRC, FunSRF and DynForest, alongside two simpler prediction approaches, to predict a relevant time-to-event outcome in each dataset, and proceeded to evaluate the predictive performance of each method according to three different metrics for survival data. Although an additional comparison with the JM approach would be interesting, JMs are not included in this benchmarking study because their estimation through state of the art statistical software for JMs (such as joineMRL and JMbayes2) repeatedly failed when attempting to include more than three to five longitudinal predictors in the JM.
@@
The remainder of this article is organized as follows: in Section 2 we briefly describe the dynamic prediction problem and the methods considered in the benchmarking study. In Section 3 we present the datasets used for the benchmarking, and provide detailed information on how each prediction method was applied to each dataset. Section 4 presents the results of the benchmarking, which are further discussed in Section 5.
2 Methods
2.1 Dynamic prediction of survival outcomes
@@
We consider the setup of a longitudinal study where $n$ subjects are enrolled, and interest lies in predicting a time-to-event outcome. Each subject $i\in\{1,2,...,n\}$ is followed from time $t=0$ (study entry / baseline visit) until they either experience the event of interest at time $t_{i}$, or they are right-censored at time $c_{i}$. We denote by $T_{i}$ and $C_{i}$ the random variables corresponding to the event and censoring times, respectively, and let $T_{i}^{*}=\min{\left(T_{i},C_{i}\right)}$. Furthermore, we denote by $t_{i}^{*}$ the observed value of $T_{i}^{*}$, and introduce the censoring indicator $\delta_{i}$ that is $1$ if for subject $i$ the event of interest is observed at $t_{i}^{*}=t_{i}$, and 00 in case of censoring, i.e., if the event has not yet been observed at $t_{i}^{*}=c_{i}$.
@@
Between times 0 and $t_{i}^{*}$, subject $i$ undergoes $m_{i}\geq 1$ visits at times $t_{i1},t_{i2},...,t_{im_{i}}$, where we assume that $t_{i1}=0$, $t_{i1}<t_{i2}<\ldots<t_{im_{i}}$ and $t_{im_{i}}<t_{i}^{*}$. At time $t_{i1}=0$, $P$ baseline covariates $x_{i}=(x_{1i},...,x_{Pi})^{T}$ are measured. Moreover, $Q$ longitudinal covariates are measured at each follow-up visit (including $t_{i1}=0$). We denote by $y_{qij}=y_{qi}(t_{ij})$ the value of the $q$-th longitudinal covariate measured on subject $i$ at time $t_{ij}$, and let $y_{ij}=y_{i}(t_{ij})=(y_{1ij},...,y_{Qij})^{T}$ be the corresponding Q-dimensional vector.
@@
Let $\mathcal{I}(t)=\{i\in\{1,2,...,n\}:t_{i}^{*}>t\}$ denote the set of individuals that are still at risk at a given time point $t$. Moreover, let $\mathcal{H}_{i}(t)=\left(x_{i},y_{i}(0),...,y_{i}(t_{ik})\right)$, where $k$ is such that $t_{ik}\leq t<t_{i,k+1}$, denote the covariate information available for a subject $i\in\mathcal{I}(t)$ up until time $t$ (in other words, $\mathcal{H}_{i}(t)$ contains the $P$ baseline covariates and all repeated measurements taken up until time $t$ of the $Q$ longitudinal covariates).
@@
At any point in time $t_{1}\in(0,t_{i}^{*})$, we may want to predict the conditional probability $P(T_{i}>t_{2}|T_{i}>t_{1})$ that individual $i\in\mathcal{I}(t_{1})$ will still be event-free at time $t_{2}>t_{1}$, given that they had not experienced the event at $t_{1}$. More specifically, given a so-called landmark time $\ell$, our goal is to estimate the conditional survival function
@@
When developing dynamic RPMs, it can often be of interest to compute predictions of survival $S_{i}(t|\ell_{k},\mathcal{H}_{i}(\ell_{k}))$ over a range of $K$ landmark times $\ell_{1},\ell_{2},...,\ell_{K}$, rather than for a single landmark $\ell$. In such situations, our notation can be easily expanded by defining $\mathcal{I}(\ell_{k})$ and $\mathcal{H}_{i}(\ell_{k})$ separately for each landmark time $\ell_{k}$.
2.2 Strict and relaxed data landmarking
Equation (1) states that the goal of dynamic prediction is to predict survival probabilities over time for all individuals who are event-free up until the landmark time $\ell$, using as predictors all covariate values measured up to time $\ell$.
In practice, often repeated measurement data gathered after $\ell$ are available when estimating a statistical model, raising the question of whether such data should be used for model estimation. Gomon et al., (2024) named “relaxed landmarking” the approach where longitudinal data gathered after the landmark are used to estimate the model on the training set, contrasting it to a “strict landmarking” method which discards any repeated measurement taken after $\ell$ in line with equation (1). They showed that using relaxed data landmarking led to a considerable worsening of the predictive performance of MFPCCox, when compared to the same model estimated using strict data landmarking.
At first sight, relaxed data landmarking may appear attractive, because it allows to use more data to estimate the model. However, this approach is problematic for two reasons: first, it uses future information (collected after $t>\ell$) to predict the conditional survival probability $P(T>t|T>\ell)$. Note that such future information is not available at the landmark time, and thus in reality it would not be possible to make use of it to make predictions at $\ell$. Furthermore, relaxed data landmarking introduces a selection bias in the modelling of the longitudinal trajectories: because individuals who survive longer are likely to have more repeated measurements taken after $\ell$ than individuals with shorter survival, the available measurements after the landmark are not a representative sample from the population of individuals who survived up until $\ell$. Whereas this is not an issue with joint models, the multi-step approaches considered in this article do not correct for this source of bias.
Despite these issues, in the literature relaxed data landmarking is often used to estimate multi-step dynamic prediction models. Based on the aforementioned remarks and the results of Gomon et al., (2024), hereafter we describe and implement the dynamic prediction models included in our comparison following a strict data landmarking approach.
2.3 Dynamic prediction methods